### v3

# Adaptive Prior Selection in Gaussian Process Bandits with Thompson Sampling

Abstract
Gaussian process (GP) bandits provide a powerful framework for performing blackbox optimization of unknown functions. The characteristics of the unknown function depend heavily on the assumed GP prior. Most work in the literature assume that this prior is known but in practice this seldom holds. Instead, practitioners often rely on maximum likelihood estimation to select the hyperparameters of the prior - which lacks theoretical guarantees. In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) that disqualifies priors with poor predictive performance, and HyperPrior GP-TS (HP-GP-TS) that utilizes a bi-level Thompson sampling scheme. We theoretically analyze the algorithms and establish upper bounds for their respective regret. In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through extensive experiments with synthetic and real-world data.
1 Introduction
In Gaussian process bandits, we consider a variant of the multi-armed bandit problem where the arms are correlated and their expected reward is sampled from a Gaussian process (GP). The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning (Turner et al., 2021), drug discovery (Hernández-Lobato et al., 2017; Pyzer-Knapp, 2018), online advertising (Nuara et al., 2018), portfolio optimization (Gonzalvez et al., 2019) and energy-efficient navigation (Sandberg et al., 2025). Most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications. Even with expert domain knowledge, selecting the exact prior to use can be a difficult task. Most practitioners tend to utilize maximum likelihood estimation (MLE) to identify suitable prior parameters. However, in a sequential decision making problem MLE is not guaranteed to recover the correct parameters.
In the literature, Wang and de Freitas (2014); Berkenkamp et al. (2019); Ziomek et al. (2024) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown. Hong et al. (2022a) introduced Thompson sampling with mixture priors (MixTS) and provided theoretical guarantees for standard and linear bandits. However, we identify apparent technical issues with their proof. More recently, Ziomek et al. (2025) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors. Their algorithm, Prior-Elimination GP-UCB (PE-GP-UCB), selects the arm and prior which provide the most optimistic upper confidence bound (UCB). If a prior generates too many incorrect predictions, then it may be eliminated. The previous work has focused on optimistic UCB methods which are known to over-explore.
In this work, we investigate the use of Thompson sampling for solving GP-bandit problems with unknown priors and we propose two algorithms. The first algorithm, Prior-Elimination GP-TS (PE-GP-TS), is an extension of PE-GP-UCB that replaces the doubly optimistic selection rule with posterior sampling and one less layer of optimism. We analyze the regret of PE-GP-TS. We obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\log T|P|\hat{\gamma}_{T}})$ (which matches that of PE-GP-UCB) plus a term (left unbounded) depending on the uncertainty of the optimal arm under the correct prior. Here, $T$ is the horizon, $|P|$ is the number of priors and $\hat{\gamma}_{T}$ is the worst-case maximum information gain. The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms.
Our analysis of HP-GP-TS yields a regret bound of order ${\mathcal{O}}(\sqrt{T\log T\bar{\gamma}_{T}})$ (where $\bar{\gamma}_{T}$ is the average maximum information gain) plus a term (left unbounded) that corresponds to the cost of learning the optimal prior.
We evaluate our methods on three sets of synthetic experiments and three experiments with real-world data. Across the experiments, our Thompson sampling based methods outperform PE-GP-UCB. Additionally, we find that the regret of HP-GP-TS does not increase with $|P|$ in our experiments. Finally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
The contributions of this work can be summarized as:
  • We propose two novel algorithms for GP-bandits with unknown prior: PE- and HP-GP-TS.
  • We theoretically analyze the regret of PE-GP-TS and HP-GP-TS.
  • We analyze the proof of the regret bound for MixTS (Hong et al., 2022a) in the linear setting, and identify technical issues.
  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve competitive performance and that the regret of HP-GP-TS does not increase with $|P|$.
2 Background and problem statement
Problem statement
We consider a sequential decision making problem where an agent repeatedly selects among a set of arms and receives a random reward whose mean depends on the selected arm and is unknown to the agent. The goal of the agent is to maximize the cumulative sum of rewards over a finite time horizon. We assume that the distribution of the means, the prior, is sampled from a set of priors, the hyperprior. An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently.
Now, let us formally state the problem. Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto[-1,1]$, $\forall p\in P$. Let $p^{*}\in P$ denote the true prior and assume the expected reward function $f:{\mathcal{X}}\mapsto{\mathbb{R}}\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ is a sample from a Gaussian process with prior $p^{*}$. Both the function $f$ and the true prior $p^{*}$ are considered unknown. We will consider two settings: In the frequentist selection setting, the prior $p^{*}\in P$ is picked arbitrarily. In the Bayesian selection setting, the prior is sampled from a hyperprior $p^{*}\sim\mathcal{P}(P)$. To simplify notation, let $P_{1}$ denote the hyperprior.
Let $T$ denote the horizon. For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $y_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d. zero-mean Gaussian noise with variance $\sigma^{2}$. The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$. In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the (potentially) stochastic selection of arms.
Gaussian processes
A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $f(x_{1}),\ldots,f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution. The probabilistic nature of GPs make them very useful for defining and solving bandit problems where the arms are correlated. Given the history $H_{t}=\{(x_{i},y_{i})\}_{i=1}^{t-1}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
$\mu_{t}(x)$ $=\mu(x)+{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}({\mathbf{y}}-{\bm{\mu}}),$ (1)
[…]
Information gain
The maximal information gain (MIG) is a measure of reduction in uncertainty of $f$ after observing the most informative data points up to a specified size. The MIG commonly occurs in regret bounds for GP bandit algorithms (Srinivas et al., 2012; Vakili et al., 2021) and its growth rate is strongly determined by the prior kernel of the GP. Hence, we will define the MIG for any fixed GP prior $p\in P$. Let ${\mathbf{y}}_{A}$ denote noisy observations of $f$ at the locations $A\subset{\mathcal{X}}$. Then, the MIG given prior $p\in P$, $\gamma_{T,p}$, is defined as
$\gamma_{T,p}:=\sup_{A\subset{\mathcal{X}},|A|\leq T}I_{p}({\mathbf{y}}_{A};f),$ (3)
where $I_{p}({\mathbf{y}}_{A};f)=H({\mathbf{y}}_{A}|p)-H({\mathbf{y}}_{A}|f,p)$ is the mutual information between ${\mathbf{y}}_{A}$ and $f$ given $p$, and $H(\cdot)$ denotes the entropy. To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{1}}[\gamma_{T,p}]$. For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ (Srinivas et al., 2012; Vakili et al., 2021).
Previous work
Plenty of previous work have proposed fully Bayesian approaches that integrate the acquisition function over the hyperposterior (Osborne et al., 2009; Benassi et al., 2011; Snoek et al., 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; De Ath et al., 2021). In contrast, HP-GP-TS optimizes a single hyperposterior sample instead of expected values over the hyperposterior. Hvarfner et al. (2023) proposed Self-Correcting Bayesian Optimization (SCoreBO) whose objective function balances reducing the uncertainty of $(x^{*},f^{*})$ and reducing the uncertainty of the true prior $p^{*}$. Notably, SCoreBO explicitly tries to identify the prior rather than integrating out the uncertainty of the prior.
Wang and de Freitas (2014) first derived regret bounds for GP bandits with unknown lengthscale for the Expected Improvement algorithm (Močkus, 1975). However, the proposed algorithm requires a lower bound on the lengthscale and the regret bound depends on the worst-case MIG. Later work by Berkenkamp et al. (2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers the lengthscale parameter. Given a sufficiently small lengthscale, the function $f$ lies within the reproducing kernel Hilbert space (RKHS) and the regular GP-UCB theory can be applied. However, A-GP-UCB lacks a stopping mechanism and will overexplore as the lengthscale continues to shrink. Recent work by Ziomek et al. (2025) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior. Unlike the work before, the regret bound of PE-GP-UCB holds for arbitrary types of hyperparameters in the GP prior. PE-GP-UCB is doubly optimistic and selects the prior and arm with the highest upper confidence bound. PE-GP-UCB tracks the cumulative prediction error made by the selected priors and eliminates priors that exceed a threshold level.
Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret (Abbasi-Yadkori et al., 2020; Pacchiano et al., 2020). Ziomek et al. (2024) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
The closest work to ours is Hong et al. (2022a). Hong et al. (2022a) studied Thompson sampling with mixture priors (MixTS) for standard bandits (with bounded rewards) and linear bandits. HP-GP-TS can be seen as an instantiation of MixTS for Gaussian process bandits. Hong et al. (2022a) provide the outline of a regret bound for a standard bandit setting and a detailed proof for the linear setting. However, we highlight some issues that we believe invalidate their analysis for the linear case in Appendix˜B. Kveton et al. (2021); Basu et al. (2021); Hong et al. (2022b); Li et al. (2024) have studied Thompson sampling based algorithms with unknown prior distributions where the agent plays multiple bandit instances. We emphasize that these methods have not been studied for GP bandits.
Figure 1: Elimination procedure of PE-GP-TS. The solid lines correspond to posterior means and the shaded regions are confidence intervals. The figure has been adapted from Ziomek et al. (2025). The dashed lines are samples from the posteriors.
3 Algorithms
As discussed by Russo and Van Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult. In addition, Thompson sampling is often observed to perform better than UCB in practice (Chapelle and Li, 2011; Wen et al., 2015; Kandasamy et al., 2018; Åkerblom et al., 2023b, a). Motivated by this, we present two algorithms for efficient prior selection based on TS.
Algorithm 1 Prior Elimination GP-TS (PE-GP-TS)
0: Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, confidence parameters $\{\beta_{t}\}_{t=1}^{T}$ and $\{\xi_{t}\}_{t=1}^{T}$.
[…]
2:for $t=1,2\ldots,T$ do
3:  Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t}$
4:  Set $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t}}\tilde{f}_{t,p}(x)$
5:  $S_{t,p_{t}}=S_{t-1,p_{t}}\cup\{t\}$ and $S_{t,p}=S_{t-1,p}$ for $p\in P\setminus\{p_{t}\}$
6:  Observe $y_{t}=f(x_{t})+\epsilon_{t}$
7:  Set $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$
8:  Set $V_{t}=\sqrt{\xi_{t}|S_{t,p_{t}}|}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})$
9:  if $\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>V_{t}$ then
10:   $P_{t+1}=P_{t}\setminus\{p_{t}\}$
11:  else
12:   $P_{t+1}=P_{t}$
3.1 Prior-Elimination with Thompson sampling
Our first algorithm is an extension of PE-GP-UCB (Ziomek et al., 2025) to be employed with Thompson sampling - instead of UCB. The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P_{t}$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t}$ where $P_{t}$ is the set of active priors. Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t}}\tilde{f}_{t,p}(x)$. Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$, PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
The elimination procedure of PE-GP-TS is illustrated in Figure˜1. Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t}$. Then, the unknown function $f$ is queried at the selected arm $x_{t}$. If the observed value differs too much from the prediction made by the selected prior, then the selected prior is eliminated. Otherwise, it remains active.
The PE-GP-TS algorithm is presented in Algorithm˜1. Similar to PE-GP-UCB, the set $S_{t,p}$ is used to store the time steps where prior $p$ was selected up to and including time $t$. When prior $p_{t}$ is selected, the prediction error $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$ between the observed and predicted value made by the prior $p_{t}$ is computed. If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line˜9. Note that at time step $t$, only the selected prior $p_{t}$ can be eliminated. As such, if a prior is very pessimistic it may never be selected and therefore will never be eliminated. Thus, the final set of active priors $P_{T}$ should be viewed as non-eliminated priors rather than necessarily being reasonable priors.
3.2 HyperPrior Thompson sampling
In our first algorithm, we removed one layer of optimism. In our second algorithm, we adopt a fully Bayesian algorithm by using a hyperposterior sampling scheme where both the prior and the mean function are sampled from their respective posteriors. By shedding the optimism over the selected prior $p_{t}$, HP-GP-TS should be able avoid costly exploration by selecting likely priors instead of optimistic ones.
The algorithm is visualized in Figure˜2 and presented in detail in Algorithm˜2. In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t}$. Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$. After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors. Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$. The algorithm can be extended to continuous priors $P$ using MCMC sampling. In comparison to SCoreBO and other fully Bayesian algorithms that compute expected values over the hyperposterior through sampling, HP-GP-TS requires only one sample from the posterior and hyperposterior – potentially reducing the computational cost significantly. The likelihood $\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ with added Gaussian noise with variance $\sigma^{2}$.
Figure 2: Overview of HP-GP-TS. The orange star represents $y_{t}$.
Algorithm 2 HyperPrior GP-TS (HP-GP-TS)
0: Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{1}$.
1:for $t=1,2\ldots,T$ do
2:  Sample $p_{t}\sim P_{t}$
3:  Sample $\tilde{f}_{t}\sim{\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$
4:  Set $x_{t}=\operatorname*{arg\,max}_{x}\tilde{f}_{t}$
5:  Observe $y_{t}=f(x_{t})+\epsilon_{t}$
6:  Set ${P}_{t+1}(p)\propto\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)\cdot P_{t}(p)$ $\triangleright$ Update hyperposterior
4 Regret analysis
In this section, we analyze the regret for the proposed algorithms. Recall from the problem statement that we consider two slightly different settings for the two algorithms. Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is selected arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{1}$.
4.1 Analysis of PE-GP-TS
Ziomek et al. (2025) structured the proof of the regret bound for PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability. Second, establishing a bound on the simple regret. Third, bounding the cumulative regret. Finally, the cumulative bound is re-expressed in terms of the worst-case MIG. For PE-GP-TS, we establish a new bound on the simple regret and then adapt the steps of Ziomek et al. to accommodate the new simple regret bound.
To bound the simple regret, we require two concentration inequalities to hold for both the posteriors and the posterior samples which we present in the following lemma.
{lemma}
[] If $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$. Then, with probability at least $1-\delta$, the following holds for all $t,x,p\in[T]\times{\mathcal{X}}\times P$:
$|f(x)-\mu_{t,p^{*}}(x)|$ $\leq\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x),$ (4)
$|\tilde{f}_{t,p}(x)-\mu_{t,p}(x)|$ $\leq\sqrt{\beta_{t}}\sigma_{t,p}(x).$ (5)
All proofs can be found in Appendix˜A. Section˜4.1 is based on Lemma 5.1 of Srinivas et al. (2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo and Van Roy (2014). Additionally, we add Equation˜5 which can be shown through the same steps and an additional union bound over $P$. Next, we state our bound for the simple regret of PE-GP-TS.
{lemma}
[] If the event of Section˜4.1 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
$f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (6)
Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$ which leads to the following regret bound:
Theorem 4.1.
Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$. If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
$R(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}\\ +2\sqrt{T\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}$ (7)
The bound of the first three terms is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t. $T$ which matches that of PE-GP-UCB. To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ (Scarlett, 2018). This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$ when considering only the first three terms. However, note that the sublinearity of ${\sum_{t\in[T]}}\sigma_{t,p^{*}}^{2}(x^{*})$ is not demonstrated.
4.2 Analysis of HP-GP-TS
We analyze the regret of HP-GP-TS using the UCB-based framework of Russo and Van Roy (2014). First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution. In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$. We decompose the regret into three terms:
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$. Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$. Together, these two terms match the Bayesian regret bound for GP-TS with a known prior. For standard GP-TS when the true prior $p^{*}$ is known, the second term has expectation zero. However for HP-GP-TS, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express the second term as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$. Hence, the second term can be seen as the cost of learning the true prior. Intuitively, if the priors are similar then the upper-confidence bounds of the true and selected prior will not differ significantly. Similarly, if the priors are sufficiently distinguishable, one would expect $P_{t}(p^{*})$ to increase quickly from a few samples. In the following lemma, we demonstrate the latter for $|P|=2$ with shared kernel functions. {lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\\ \times\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$. In Equation˜8, $\bm{\mu}$} is the difference in mean between the two priors for the given sequence of arms and ${\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}$ is the whitened difference vector. As the norm of the whitened difference vector increases, $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, the sum of the two rightmost terms in Equation˜9 goes to zero from below. The norm $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, and consequently the posterior probability of the true prior $P_{t+1}(p^{*})$, is maximized when arms with large difference in prior mean (relative to their variance) and with low correlation are selected. Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting.
Next, we state the Bayesian regret bound for HP-GP-TS.
Theorem 4.2.
If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$, then the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]\\ +\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors. This is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$ when considering only the first and third terms. We have not shown that ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is sublinear. As noted before, this sum corresponds to the cost of learning the true prior. Conceptually, we expect the terms in the sum to decrease over time as either $p_{t}$ or $x_{t}$ converge to $p^{*}$ or $x^{*}$ respectively. Empirically, we observe this in our synthetic experiments, see Figure˜13 in Appendix˜F.
4.3 Comparison to MixTS
Hong et al. (2022a) study MixTS, a Thompson sampling algorithm that assumes the prior is a mixture distribution, for standard and linear bandits. Using our GP-notation, their analysis starts from the decomposition
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}[\underbrace{f(x^{*})-\mu_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{\mu_{t,p_{t}}(x_{t})-f(x_{t})}_{(2)}].$ (11)
The first term is bounded using standard UCB techniques. For the second term, Hong et al. (2022a) introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ with high probability. Their key idea is to introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ holds with high probability, similar to Ziomek et al. (2025). In our GP setting, their excess reward can be defined as
$G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(a_{s})-\sqrt{\beta_{t}}\sigma_{s,p}(a_{s})-y_{s})$ (12)
and the prior confidence set $C_{t}=\{p\in P:G_{t}(p)\leq 2\sigma\sqrt{N_{t}(p)\log T}\}$ where $N_{t}(p)$ is the number of times prior $p$ has been selected before time $t$. Unlike the definition of Ziomek et al. (2025), priors can return to the confidence set after previously being removed from it. When bounding the second term in their decomposition for a standard bandit with bounded rewards, Hong et al. (2022a) condition on $p_{t}\in C_{t}$. If $p_{t}\in C_{t}$, the second term can be bounded by all the excess $\sum_{p\in P}G_{T+1}(p)$. If $p_{t}\notin C_{t}$ and the rewards are in $[0,1]$, then the second term is bounded by the probability of $p^{*}\in C_{t}$. For the linear setting with unbounded rewards, the proof requires conditioning on the linear parameter vector $\theta^{*}$ to lie close to its prior mean. Under this additional event $E_{0}$, the distribution of the true parameter vector $\theta^{*}$, the true prior and the optimal arm can shift. But, conditioned on the history $H_{t}$, MixTS is unaffected by $E_{0}$ at time step $t$. However, Hong et al. (2022a) use that $x^{*},p^{*}|H_{t},E_{0}\overset{d}{=}x_{t},p_{t}|H_{t},E_{0}$ without proof, potentially invalidating Theorem 1 of Hong et al. (2022a). We describe this and other technical issues in detail in Appendix˜B.
Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The final regret for PE-GP-UCB is 114 and 389 in the lengthscale and subspace experiments, and 181 for SCoreBO in the lengthscale experiment. Errorbars correspond to $\pm 1$ standard error.
5 Experiments
In this section, we describe our experiments based on synthetic and real-world data.
Synthetic experiments
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$. For the second setup, 8 priors use the RBF kernel with different lengthscales equidistantly spaced between $1/2$ and $4$. For the third setup, the total dimensions $d=16$ but each of the 5 priors $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions. The 4 subdimensions are designed such that the priors are equally difficult to distinguish. All priors use the RBF kernel with lengthscale $\ell=8$. For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB, SCoreBO, fully Bayesian Expected Improvement (EEI) and Maximum A Posteriori (MAP) GP-TS. MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation.. In addition, we compare against the oracle variants of PE-GP-TS and PE-GP-UCB that are only given the true prior: $P_{1}=\{p^{*}\}$.
The cumulative regret for the three synthetic experiments is shown in Figure˜3 and the final regret is shown in Table˜1 in Appendix˜F. Across all three experiments, we observe that HP-GP-TS and EEI has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment. The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances. SCoreBO has the highest regret in the kernel and lengthscale experiment but has more comparable performance in the subspace experiment. In Figure˜10 in Appendix˜F, we report the regret for the two most competitive methods, HP-GP-TS and EEI, with an extended horizon $T=1500$ where we observe that HP-GP-TS yields noticeably lower regret.
The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Figure˜4. The PE-methods eliminate at most one prior on average. In contrast, the final hyperposterior entropy across all algorithms is equivalent to 70-99% of the probability mass being assigned to one prior showing that the hyperposterior adapts more effectively. Across the experiments, SCoreBO has the lowest hyperposterior entropy followed by HP-GP-TS and EEI has the highest (except for the kernel experiment), see Figures˜11 and 12 in Appendix˜F. Thus, HP-GP-TS has similar regret to EEI but lower hyperposterior entropy.
In Figure˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices. PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds. The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy" of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.9\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data. See Figure˜14 in Appendix˜F for confusion matrices in the lengthscale and subspace experiments.
Figure 4: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS (left). Mean entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS (right). The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
Figure 5: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ for the kernel experiment. Row-wise normalized to 100%.
Figure 6: Total regret for the lengthscale and subspace experiments as $|P|$ increases.
Figure 7: Cumulative regret on the real-world data experiments. Errorbars correspond to $\pm 1$ standard error. The average final regret for SCoreBO and PE-GP-UCB is 861 and 506 on PNW.
Scaling $|P|$
We perform two experiments to understand how the regret of our algorithms scale with the number of priors. In both experiments, the average difficulty of the problem is kept constant such that the regret of the oracle models is constant. In the first experiment, we increase the discretization of the lengthscale values. The lengthscales are equidistantly spaced in $[0.5,4]$ with $|P|\in\{8,16,32,64,128\}$. As $|P|$ increases, the difference between similar priors is reduced. In the second experiment, we increase the number of priors in the subspace experiment from 5 up to 16. Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different. The total regret as the number of priors increases is shown in Figure˜6. For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm, likely due to the increased redundancy in the priors. The one exception is SCoreBO, whose regret increases as $|P|$ is increased to 32 but levels off beyond that. In the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle. The regret of EEI drops initially but is otherwise constant and SCoreBO also has constant regret.
Real-world data
We perform three experiments with real-world data from the Intel Berkeley dataset (Madden and others, 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann and Bretherton, 1999, 2000). Each dataset contains measurements from a set of sensors over time. We split each dataset into a training and test set where the test set contains the last third of the data. Hence, the distribution of the test data may have shifted slightly from the training data. The training sets are split further into separate buckets to define our priors. For each bucket $p$, we compute the empirical mean $\hat{\mu}_{p}$ and covariance $\hat{\Sigma}_{p}$ which defines the prior ${\mathcal{GP}}(\hat{\mu}_{p},\hat{\Sigma}_{p})$. The buckets in the Intel data corresponds to the 12 days in the training dataset. For the PeMS data, each hour between 06:00 and 13:00 defines one bucket, giving 7 priors. For the daily precipitation data, each month in the year constitutes a bucket, yielding 12 priors. When running the experiments, we select a measurement of all sensors from the test data uniformly at random. The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipitation respectively for the three datasets. When the algorithms select an arm to evaluate, we add Gaussian noise with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al. (2012); Bogunovic et al. (2016). See Appendix˜E for details about the experimental setup.
The cumulative regret for the experiments with real-world data is presented in Figure˜7. Across these experiments, HP-GP-TS has either the lowest regret or is within 1 standard error of the algorithm with the lowest regret. SCoreBO has significantly higher regret than all other methods. Notably for the PeMS data, PE-GP-TS has the lowest average regret whereas MAP GP-TS and EEI perform worse compared to the other experiments.
The number of priors remaining in $|P_{t}|$ and the hyperposterior entropy for the real-world data experiments is shown in Figures˜11 and 12 in Appendix˜F. Similar to the synthetic experiments, on average, the prior elimination methods eliminate less than 1 prior at best and no priors (across all 500 seeds) at worst. In contrast, the hyperposterior of HP-GP-TS concentrates to the equivalent of 60-80% of the probability mass to one prior. The relative standing in terms of reduced hyperposterior uncertainty between SCoreBO, HP-GP-TS and EEI remains consistent across all the experiments. SCoreBO reduces the hyperposterior uncertainty the most at the cost of significantly higher regret whereas HP-GP-TS provides a better balance between low regret and low hyperposterior uncertainty.
6 Conclusion
In this paper, we have proposed two algorithms for adaptive prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data. We find that they both select the true prior more often and obtain lower or comparable regret than previous work due to lowering the amount of optimistic exploration.
Acknowledgments
The work of Jack Sandberg and Morteza Haghir Chehreghani was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Proofs
In the following section, we state and prove the results shown in the main text.
A.1 PE-GP-TS
First, we state and prove concentration inequalities for $f(x)$ and $\tilde{f}_{t,p}(x)$.
See 4.1
Proof.
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Equation˜5 we must take an additional union bound over $p\in P$.
Fix $t,x,p\in[T]\times{\mathcal{X}}\times P$. Given the history $H_{t}$, $\tilde{f}_{t,p}(x)\sim{\mathcal{N}}(\mu_{t,p}(x),\sigma^{2}_{t,p}(x))$. Using that $\mathbb{P}(Z>c)\leq 1/2e^{-c^{2}/2}$ for $Z\sim{\mathcal{N}}(0,1)$, we get that
$\mathbb{P}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p}(x)}{\sigma_{t,p}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (13)
$=\frac{3\delta}{|{\mathcal{X}}||P|\pi^{2}t^{2}}$ (14)
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Equation˜5 holds w.p. at least $1-\delta/2$. By the same reasoning and skipping the union bound over $P$, Equation˜4 holds w.p. at least $1-\delta/2$. Thus, both events hold w.p. at least $1-\delta$. ∎
Next, we state three lemmas from Ziomek et al. (2025) that are used in the proof of our regret bound.
{lemma}
(Lemma 5.1 of Ziomek et al. (2025)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
$\left|\sum_{i\in S_{t,p}}\epsilon_{i}\right|\leq\sqrt{\xi_{t}|S_{t,p}|}\quad\forall t,p\in[T]\times P.$ (15)
{lemma}
(Lemma 5.2 of Ziomek et al. (2025)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section˜4.1 holds, then
$\sup_{x\in{\mathcal{X}}}|f(x)|\leq B_{p^{*}}.$ (16)
{lemma}
(Lemma 5.3 of Ziomek et al. (2025)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
Then, we state and prove the new simple regret bound for PE-GP-TS.
See 4.1
Proof.
First, we upper bound $f(x^{*})$ as follows
[…]
$\leq\tilde{f}_{t,p_{t}}(x_{t})+2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*}).$ $(\text{TS selection rule})$ (19)
Then, we lower bound $f(x_{t})$
$f(x_{t})$ $=\mu_{t,p_{t}}(x_{t})+\eta_{t}-\epsilon_{t}$ $(\text{Def. of }\eta_{t})$ (20)
[…]
Combining, Equations˜19 and 21 we obtain
$f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (22)
Finally, we state and prove the cumulative regret bound for PE-GP-TS.
See 4.1
Proof.
First, we show that the true prior $p^{*}$ is never rejected if Sections˜4.1 and A.1 hold.
$\left|\sum_{i\in S_{t,p^{*}}}\eta_{i}\right|$ $=\left|\sum_{i\in S_{t,p^{*}}}(y_{i}-f(x_{i})+f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ (23)
$\leq\left|\sum_{i\in S_{t,p^{*}}}\epsilon_{i}\right|+\sum_{i\in S_{t,p^{*}}}\left|f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ $(\text{Triangle ineq.)}$ (24)
[…]
Next, we bound the cumulative regret. To establish a bound on the cumulative regret, we must separate out the rounds where priors are eliminated. Hence, define the set of critical iterations as
$\mathcal{C}=\left\{t\in[T]:\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}S_{t,p_{t}}}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})\right\}.$ (26)
Using Equations˜15 and 22, we can bound the cumulative regret as follows:
${\text{BR}}(T)$ $=\sum_{t\in\mathcal{C}}{\text{BR}}_{t}+\sum_{t\notin\mathcal{C}}{\text{BR}}_{t}$ (27)
$\leq 2|P|B_{p^{*}}+\sum_{t\notin\mathcal{C}}2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}(\epsilon_{t}-\eta_{t}).$ (28)
where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin\mathcal{C}$, line˜9 in Algorithm˜1 evaluates to false and hence
$\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}-\eta_{t}\leq\sum_{p\in P}\sqrt{\xi_{T}|S_{T,p}|}+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p}(x_{t}).$ (29)
Additionally, using Section˜A.1, we can bound the Gaussian noise:
$\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}\epsilon_{t}$ $\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}\setminus\mathcal{C}}\epsilon_{t}\right|\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}}\epsilon_{t}\right|$ (30)
[…]
$\leq\sqrt{\xi_{T}|P|T}$ $(\text{Cauchy-Schwarz})$ (32)
Combining the above, the cumulative regret is bounded by
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t}).$ (33)
Finally, applying Equation˜16, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{T\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (34)
A.2 UCB-analysis of HP-GP-TS
Next, we state and prove our UCB-based regret bound for HP-GP-TS.
See 4.2
Proof.
To begin, we note that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ and $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since both $x_{t}$ and $p_{t}$ are sampled from their respective posteriors. Let $U_{t,p}(x):=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$. Then, we start decomposing the instant regret into two terms:
${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-f(x_{t})\right]$ (35)
$={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})+U_{t,p^{*}}(x^{*})$ (36)
$\qquad\qquad-U_{t,p_{t}}(x^{*})+U_{t,p^{*}}(x_{t})-f(x_{t})]\hskip 60.00009pt(x^{*},p_{t}|H_{t}\overset{d}{=}x_{t},p^{*}|H_{t})$ (37)
$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-U_{t,p^{*}}(x^{*})\right]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-U_{t,p^{*}}(x_{t})\right]}_{(2)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x_{t})-f(x_{t})\right]}_{(3)}$ (38)
We begin by bounding term $(1)$,
$(1)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-\mu_{t,p^{*}}(x^{*})-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})\right]$ (39)
$\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\left[f(x^{*})-\mu_{t,p^{*}}(x^{*})-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})\right]_{+}\right]$ $\left([\cdot]_{+}:=\max(\cdot,0)\right)$ (40)
$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\right]$ $(x^{*}\in{\mathcal{X}},[\cdot]_{+}\geq 0)$ (41)
$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[{\mathbb{E}}_{t}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\big|\,p^{*},H_{t}\right]\right]$ $(\text{Tower rule})$ (42)
Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]\leq\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$. In our case, note that $f(x)|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}}(x),\sigma^{2}_{t,p^{*}}(x))$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$. Hence,
$(1)$ $\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[\frac{\sigma_{t,p^{*}}(x)}{\sqrt{2\pi}}\exp\left(\frac{-\beta_{t}}{2}\right)\right]$ (43)
$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{\beta_{t}}{2}\right)\right]$ $(\sigma_{t,p^{*}}(x)\leq\sigma_{0,p^{*}}(x)\leq 1)$ (44)
$={\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}\frac{1}{\sqrt{2\pi}}\exp(-\beta_{t}/2)$ (45)
$\leq{\sum_{t\in[T]}}\frac{1}{t^{2}}\leq\frac{\pi^{2}}{6}.$ $(\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi}))$ (46)
Next, we bound $(3)$ as follows:
$(3)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x_{t})-f(x_{t})\right]$ (47)
$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p^{*}}(x_{t})-f(x_{t})|H_{t}\right]\right]$ $(\text{Tower rule})$ (48)
$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p^{*}}(x_{t})-\mu_{t,p^{*}}(x_{t})|H_{t}\right]\right]$ $({\mathbb{E}}[f(x_{t})|H_{t}]={\mathbb{E}}[\mu_{t,p^{*}}(x_{t})|H_{t}])$ (49)
$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}[{\mathbb{E}}_{t}[\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})|H_{t}]]$ $(U_{t,p^{*}}(\cdot)=\mu_{t,p^{*}}(\cdot)+\sqrt{\beta_{t}}\sigma_{t,p^{*}}(\cdot))$ (50)
$={\sum_{t\in[T]}}{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})]$ (51)
Continuing,
$(3)$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})\right]$ (52)
$\leq{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\beta_{t}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\text{Cauchy-Schwarz})$ (53)
$=\sqrt{{\sum_{t\in[T]}}\beta_{t}}{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\beta_{t}\text{ deterministic})$ (54)
$\leq\sqrt{{\sum_{t\in[T]}}\beta_{t}}\sqrt{{\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\right]}$ $(\text{Jensen's inequality})$ (55)
$\leq\sqrt{\beta_{T}T}\sqrt{{\mathbb{E}}_{p^{*}}\left[{\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\,\bigg|\,p^{*}\right]\right]}$ $(\beta_{t}\text{ increasing})$ (56)
[…]
$\leq\sqrt{\beta_{T}T}\sqrt{C\bar{\gamma}_{T}}$ (58)
Combining the bounds for (1) and (3), we obtain the desired result
${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-U_{t,p^{*}}(x_{t})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (59)
A.3 Analysis of expected posterior probabilities
See 8
Proof.
Fix $p^{*}=p$ and let $\tilde{p}$ denote the incorrect prior. Without loss of generality, assume $\mu_{1,p}(x)=0$ $\forall x\in{\mathcal{X}}$ such that the two priors are: ${\mathcal{GP}}_{p}(0,k)$ and ${\mathcal{GP}}_{\tilde{p}}(\mu,k)$. For the observed rewards ${\mathbf{y}}\in{\mathbb{R}}^{t}$, we use $N_{p}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};\mathbf{0},{\mathbf{K}}+\sigma^{2}I)$ and $N_{\tilde{p}}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};{\bm{\mu}}(x_{1:t}),{\mathbf{K}}+\sigma^{2}I)$ to denote the likelihood (or multivariate Gaussian density) under the priors $p$ and $\tilde{p}$ where ${\bm{\mu}}(x_{1:t})\in{\mathbb{R}}^{t}$ and ${\mathbf{K}}\in{\mathbb{R}}^{t\times t}$ s.t. $({\bm{\mu}})_{i}=\mu(x_{i})$ and $({\mathbf{K}})_{i,j}=k(x_{i},x_{j})$. To simplify the notation, let ${\bm{\mu}}:={\bm{\mu}}(x_{1:t})$ and ${\bm{\Sigma}}=K+\sigma^{2}I$ s.t. $N_{p}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};\bf{0},{\bm{\Sigma}})$ and $N_{\tilde{p}}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};{\bm{\mu}},{\bm{\Sigma}})$.
By Bayes’ theorem, the posterior probability of $p$ satisfies:
$P_{t+1}(p)$ $=\frac{N_{p}({\mathbf{y}})P_{0}(p)}{N_{p}({\mathbf{y}})P_{0}(p)+N_{\tilde{p}}({\mathbf{y}})P_{0}(\tilde{p})}=\frac{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)}{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)+(1-P_{0}(p))}.$ (60)
Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}-{\bm{\mu}})^{T}{\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$. Since we wish to lower bound the expectation of $P_{t+1}(p)$, we will utilize the lower bounds $q(x)\geq ce^{x}$ $\forall x\leq 0$ and $q(x)\geq 1-\frac{1-c}{c}e^{-x}$ $\forall x\in{\mathbb{R}}$. The first lower bound will be applied to the region $H:=\{{\mathbf{y}}\big|Q({\mathbf{y}})\leq 0\}$ whilst the second is applied to the complement $H^{c}$:
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $=\frac{1}{\sqrt{(2\pi)^{t}}}\int_{{\mathbb{R}}^{t}}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\frac{\exp(Q({\mathbf{y}}))c}{\exp(Q({\mathbf{y}})c+1-c}\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}\right)d{\mathbf{y}}$ (61)
$\geq\frac{1}{\sqrt{(2\pi)^{t}}}\Bigg(c\underbrace{\int_{H}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\exp(Q({\mathbf{y}}))\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}})\right)d{\mathbf{y}}}_{I_{1}:=}$ (62)
$+\underbrace{\int_{H^{c}}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\left(1-\frac{1-c}{c}\exp(-Q({\mathbf{y}}))\right)\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}})\right)d{\mathbf{y}}}_{I_{2}:=}\Bigg).$ (63)
Let $\bm{\Sigma}^{\frac{1}{2}}$} be the positive definite and symmetric square root of $\bm{\Sigma}$} s.t. ${\bm{\Sigma}}={\bm{\Sigma}^{\frac{1}{2}}}{\bm{\Sigma}^{\frac{1}{2}}}$. Next, we apply the change of variables ${\mathbf{y}}={\bm{\Sigma}^{\frac{1}{2}}}(v\bar{{\bm{\mu}}}+{\bm{w}})$ s.t. ${\bm{w}}\perp\bar{{\bm{\mu}}}$ and $\bar{{\bm{\mu}}}:={\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}/\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$. Note that the Jacobian of this variable change is $\sqrt{\det{\bm{\Sigma}}}$ and
${\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}=(v\bar{{\bm{\mu}}}+{\bm{w}})^{T}{\bm{\Sigma}^{\frac{1}{2}}}{\bm{\Sigma}}^{-1}{\bm{\Sigma}^{\frac{1}{2}}}(v\bar{{\bm{\mu}}}+{\bm{w}})=\|v\bar{{\bm{\mu}}}+{\bm{w}}\|^{2}=v^{2}+\|{\bm{w}}\|^{2},$ (64)
$({\mathbf{y}}-{\bm{\mu}})^{T}{\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})$ $=((v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)\bar{{\bm{\mu}}}+{\bm{w}})^{T}((v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)\bar{{\bm{\mu}}}+{\bm{w}})$ (65)
$=(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}+\|{\bm{w}}\|^{2},$ (66)
$Q({\mathbf{y}})=-\frac{1}{2}\left(v^{2}+\lVert{\bm{w}}\rVert^{2}-(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}-\lVert{\bm{w}}\rVert^{2}\right)=-\frac{1}{2}\left(2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}\right).$ (67)
The region $H=\{Q({\mathbf{y}})\leq 0\}$ can then be expressed as $H=\left\{(v,{\bm{w}})\big|v\geq\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2,{\bm{w}}\perp{\bm{\mu}}\right\}$ and $H^{c}=\{(v,{\bm{w}})\big|v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2,{\bm{w}}\perp{\bm{\mu}}\}$. Next, we apply the change of variables to $I_{1}$ and decompose the integral into two parts:
$I_{1}$ $=\int_{H}\exp\left(-\frac{1}{2}\left(-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+v^{2}+\lVert{\bm{w}}\rVert^{2}\right)\right)dvd{\bm{w}}$ (68)
$=\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\int_{v\geq\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{1}{2}(v+\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{\lVert{\bm{w}}\rVert^{2}}{2}\right)d{\bm{w}}$ (69)
$=\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\cdot\sqrt{2\pi}\left(1-\Phi\left(\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)\right)\cdot\sqrt{(2\pi)^{t-1}}$ (70)
$=\sqrt{(2\pi)^{t}}\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\Phi\left(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)\qquad\qquad\qquad\qquad\left(\Phi(-x)=1-\Phi(x)\right)$ (71)
where $\Phi(\cdot)$ is the CDF of the unit-Gaussian. Similarly, we apply the change of variables to $I_{2}$ and decompose the integral into four parts:
$I_{2}$ $=\int_{H^{c}}\Bigg(\exp\left(-\frac{1}{2}\left(v^{2}+\lVert{\bm{w}}\rVert^{2}\right)\right)$ (72)
$\qquad\qquad-\frac{1-c}{c}\exp\left(-\frac{1}{2}(-2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+v^{2}+\lVert{\bm{w}}\rVert^{2})\right)\Bigg)dvd{\bm{w}}$ (73)
$=\int_{v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{v^{2}}{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{w^{2}}{2}\right)d{\bm{w}}$ (74)
$\quad-\frac{1-c}{c}\int_{v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{1}{2}\left(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)^{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{2}\right)d{\bm{w}}$ (75)
$=\sqrt{2\pi}\Phi\left(\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\cdot\sqrt{(2\pi)^{t-1}}-\frac{1-c}{c}\sqrt{2\pi}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\cdot\sqrt{(2\pi)^{t-1}}$ (76)
$=\sqrt{(2\pi)^{t}}\left(1-\frac{1}{c}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\right).\qquad\qquad\qquad\qquad\left(\Phi(-x)=1-\Phi(x)\right)$ (77)
Combining the results for $I_{1}$ and $I_{2}$, we get that
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\left(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)-\frac{1}{P_{0}(p)}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)$ (78)
Note that a sharp bound of Equation˜78 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz and Stegun, 1972):
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{3\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{9\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+16}}$ (79)
$\quad-\frac{1}{P_{0}(p)}\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+32\pi^{-1}}}.$ (80)
Appendix B Technical issues with MixTS regret bound in the linear setting
Theorem 1 in Hong et al. (2022a) provides a regret bound for MixTS in the linear setting. The linear setting assumes that the true parameter $\theta^{*}\sim{\mathcal{N}}(\theta_{0,S_{*}},\Sigma_{0,S_{*}})$ where the latent state $S_{*}$ is sampled from a discrete mixture prior $P_{0}$. The proof of Theorem 1 in Hong et al. (2022a) contains a few non-obvious steps that seem difficult to motivate. We use the notation of Hong et al. (2022a).
  1. 1.
    Eq. (9) in Hong et al. (2022a) uses the TS property that the true prior and optimal arm is equal in distribution to the selected prior and selected arm given the history: $A_{t,*}^{\top}\bar{\theta}_{t,S_{*}}|H_{t}\overset{d}{=}A_{t}^{\top}\bar{\theta}_{t,S_{t}}|H_{t}$ (equivalent to $\mu_{t,p^{*}}(x^{*})|H_{t}\overset{d}{=}\mu_{t,p_{t}}(x_{t})|H_{t}$ in our notation). However, Eq. (9) additionally conditions on the event $E_{0}=\{\|\theta_{*}-\theta_{0,S_{*}}\|_{\Sigma_{0,S_{*}}^{-1}}\leq\sqrt{2d\log(dn)}\}$ where $\theta_{*}$ lies close to its prior mean. The TS property does not hold under this event since it modifies the distribution of the linear parameter $\theta^{*}$ but not the sampled parameters $\theta_{t}$, thus changing the distribution of the optimal arm $A_{t,*}$ but not the selected arm $A_{t}$. Consider the example in Figure˜8, if $E_{0}$ holds then $\theta_{*}$ lies in the blue regions and thus $a_{2}$ is optimal w.p. 0. If $E_{0}^{c}$ holds, then $a_{2}$ is optimal with a non-zero probability. However, MixTS is oblivious to $E_{0}$ given the history and thus $A_{t,*}|H_{t},E_{0}\not\overset{d}{=}A_{t}|H_{t}$. This counterexample illustrates the overall idea but we have not validated that the scale of the arms and the blue regions are feasible.
  2. 2.
    Five lines above Eq. (9) in Hong et al. (2022a), it is stated that the regret is upper-bounded by a constant $M$ whenever $E_{0}$ occurs. However, from Eq (9) to the first term in step 3 of their analysis (page 15), the bound of $M$ is applied implicitly to $A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}$ without motivation. For the setting with bounded rewards, then $\bar{\theta}_{t,S_{t}}$ is also bounded but for Gaussian rewards $\bar{\theta}_{t,S_{t}}$ can be unbounded.
  3. 3.
    The second term in Eq. (9) contains the indicator function $\mathbf{1}\{E_{0}\}$: $\mathbb{E}[(A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*})\mathbf{1}\{E_{0}\}]$. In step 3 (page 15), this indicator function is dropped without motivation: $\mathbb{E}[\langle A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}\rangle_{M}]$ where $\langle\cdot\rangle_{M}=\min(\cdot,M)$ for the bound $M$. If the expression inside is non-negative w.p. 1, then this step would be valid but this is not the case.
  4. 4.
    From our understanding, the final equation on page 15 adds and subtracts the confidence bound and adds a zero-mean Gaussian inside a minimum. However, adding a zero-mean Gaussian inside a minimum reduces the expectation but the analysis seems to assume that it would increase the expectation. I.e. it is seemingly assumed that $\mathbb{E}[\min(M,X)]\leq\mathbb{E}[\min(M,X+\epsilon_{t})$ for a constant M and random variable $X$. However, the reverse inequality is true.
Theorem 1 of Hong et al. (2022a) provides a novel proof technique that could potentially be applied to HP-GP-TS. But the technical issues presented above present a few challenges.
Figure 8: Potential counterexample to Eq (9) in Hong et al. (2022a). The blue regions represent when the event $E_{0}$ holds and the black dots represent the two arms $a_{1}$ and $a_{2}$.
Appendix C Extending the analysis of MixTS to HP-GP-TS
In this section, we explore how the analysis of Hong et al. (2022a) can be applied to the Gaussian process setting considered in this paper. The core idea of the analysis of Hong et al. (2022a) is to introduce confidence sets for the priors (or latent variables in their terminology). The confidence intervals closely resemble the prior elimination criteria introduced by Ziomek et al. (2025). In our notation, the equivalent confidence set $C_{t}$ of Hong et al. (2022a) is defined as
$C_{t}=\left\{p\in P:\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})\leq 2\sigma\sqrt{N_{t}(p)\log T}\right\}.$ (81)
We define $G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})$. The prior elimination criteria for PE-GP-UCB and PE-GP-TS can be expressed as a confidence set in the following way:
$\tilde{C}_{t}=\Bigg\{p\in P:\forall s<t\quad\Bigg|\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}y_{i}-\mu_{i,p}(x_{i})\Bigg|\leq\sqrt{2\sigma^{2}N_{t}(p)\log(t^{2}|P|\pi^{2}/3)}+\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}\sqrt{\beta_{i}}\sigma_{i,p}(x_{i})\Bigg\}.$ (82)
The main difference between $C_{t}$ and $\tilde{C}_{t}$ is that $\tilde{C}_{t}$ is monotonically decreasing in cardinality since eliminated priors cannot be selected by PE-GP-UCB and -TS. Additionally, $\tilde{C}_{t}$ uses a symmetric criteria and the confidence parameters in $C_{t}$ use the final time step $T$ rather than the current time step $t$.
To apply the analysis of Hong et al. (2022a) directly to a GP-setting would require resolving the issues discussed in Appendix˜B. To circumvent some of these issues, we attempt to analyze the regret of HP-GP-TS by directly applying a UCB-decomposition and manipulate the decomposition such an expression similar to $G_{t}(p_{t})$ appears.
Recall, the regular UCB-decomposition to the Bayesian cumulative regret:
${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-f(x_{t})]$ (83)
$={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})+U_{t,p^{*}}(x^{*})-f(x_{t})]$ (84)
$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p_{t}}(x_{t})-f(x_{t})]}_{(2)}\quad(p_{t},x_{t}|H_{t}\overset{d}{=}p^{*},x^{*}|H_{t})$ (85)
By the same steps as in the proof of Theorem˜4.2, $(1)\leq\pi^{2}/6$. Hence, we focus on the second term and condition on whether the selected prior lies in the confidence set:
$(2)={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]+{\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\notin C_{t}\}].$ (86)
To bound the right-term above, we would like $(U_{t,p_{t}}(x_{t})-f(x_{t}))$ to be bounded such that it is sufficient to analyze the probability $\mathds{1}\{p_{t}\notin C_{t}\}$. However, we have not identified a method of doing so. We continue to study the left term. Since the noise $\epsilon_{t}$ is independent of $\mathds{1}\{p_{t}\in C_{t}\}$ and has zero mean, we can add it inside the expectation and add-and-subtract the confidence radius:
${\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]$ (87)
$={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (88)
$={\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (89)
$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (90)
$\leq\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]}_{(3)}$ (91)
$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})]$ $(\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\geq 0)$ (92)
Note that the sum of $\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})$ can be bounded in terms of the worst-case MIG $\hat{\gamma}_{T}$} by Equation˜16. To bound (3), following the argument of Hong et al. (2022a), we note that for each prior $p\in P$ there must be some last time $t^{\prime}(p)\leq T$ that $p$ was in the confidence set and chosen by the algorithm: $p\in C_{t^{\prime}}$ and $p=p_{t^{\prime}}$. Since $p\in C_{t^{\prime}}$, then $G_{t^{\prime}}(p)\leq 2\sigma\sqrt{N_{t^{\prime}}(p)\log T}$. Therefore, for each prior $p$, we can split the sum into $G_{t^{\prime}}(p)$ and the final term at time $t^{\prime}$:
$(3)$ $\leq{\mathbb{E}}\bigg[\sum_{p\in P}(G_{t^{\prime}}(p)+(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg]$ (93)
$\leq{\mathbb{E}}\bigg[\sum_{p\in P}2\sigma\sqrt{N_{t^{\prime}}(p)\log T}\bigg]$ (94)
$\quad+{\mathbb{E}}\bigg[\sum_{p\in P}(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg].$ (95)
The first sum above can be bounded using that $\sum_{p\in P}N_{T}(p)=T$ and Jensen’s inequality. The major challenge is to show that the final term is bounded. This requires creating an event s.t. the terms inside are bounded with high probability. However, as the issues in Appendix˜B highlight, if the event is in introduced at an earlier step then it must not invalidate the subsequent manipulations.
Appendix D Description of kernels
[…]
Appendix E Additional experimental details
In this section, we provide some additional details about the experiments.
For the kernel experiment, all kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors is zero everywhere. For the subspace experiment, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$. Dimensions larger than $5$ are wrapped around 1, i.e. $((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize. The prior elimination methods use $\delta=0.05$ across all experiments, including the oracle methods. During every iteration $t$, SCoreBO samples $M$ priors from the hyperposterior $P_{t}$ and samples $N$ optimizers $x^{*},f^{*}$ for each prior sampled through posterior sampling. We use $M=16$ and $N=12$ for SCoreBO in all experiments. We use the implementation of the SCoreBO acquisition function in BoTorch (Community) (Balandat et al., 2020). To make the implementation fast with GPUs, we set linear_operator.settings.stable_qr_cpu_threshold to 8 in order to avoid QR-factorization being performed on CPU (Gardner et al., 2018; Pleiss et al., 2022, 2025). To avoid out of memory issues, we replace the default torch.matmul in DefaultPredictionStrategy._exact_predictive_covar_inv_quad_form_root (from gpytorch.models.exact_prediction_strategies) with an equivalent torch.einsum (Gardner et al., 2018). Since the priors in our experiments are discrete, we compute the hyperposterior exactly and sample from it directly.
For all the real-world datasets, sensors containing any null measurements are filtered out.
The Intel Berkeley dataset consists of measurements from 46 temperature sensors across 19 days. The training set consists of the first 12 days of measurements and the remaining 7 days constitute the test set. The noise variance is set to $\sigma^{2}=0.7^{2}$.
The PeMS data consists of measurements from 211 sensors along the I-880 highway from all of 2023. The goal is to find the sensors with low speeds to identify congestions. We use the 5-min averages provided by PeMS. Data between 2023-01-01 and 2023-09-01 is put into the training set whilst the data until 2023-12-31 is put into the test set. The noise variance is set to $\sigma^{2}=2.25^{2}$.
The PNW precipitation data consists of daily precipitation data from 1949 to 1994 across 167 $50\times 50$ km regions in the Pacific Northwest. The goal is to find the region with the highest precipitation for any given day. The training data consists of the measurements made prior to 1980 and the test data consists of the measurements between 1980 and 1994. The original data is stated to be given in mm/day however the data seems to be off by a factor of 10. We rescale the data to a log-scale using $\log(\cdot/10+0.1)$, similar to Krause et al. (2008). The noise variance is set to $\sigma^{2}=0.41^{2}$.
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C on this instance, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Figure˜9. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be an indication that this particular sensor was starting to fail earlier than others.
Figure 9: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures.
Appendix F Additional experimental results
In this section, we provide some additional experimental results.
Table 1: Average total regret and $\pm$ 1 standard error for the synthetic and real-world data experiments. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Synthetic Real-world data
Kernel Lengthscale Subspace Intel PeMS PNW Precip.
MAP GP-TS 84.3 $\pm$ 8.4 30.2 $\pm$ 1.2 87.2 $\pm$ 1.0 73.8 $\pm$ 7.7 1635.0 $\pm$ 129.3 178.4 $\pm$ 6.9
HP-GP-TS 39.2 $\pm$ 1.4 31.4 $\pm$ 1.0 88.3 $\pm$ 0.9 54.1 $\pm$ 3.0 1327.8 $\pm$ 107.9 167.7 $\pm$ 5.2
PE-GP-TS 62.0 $\pm$ 0.6 61.8 $\pm$ 0.5 177.1 $\pm$ 1.4 106.5 $\pm$ 2.1 1214.2 $\pm$ 81.5 200.9 $\pm$ 4.0
PE-GP-UCB 121.6 $\pm$ 1.2 114.2 $\pm$ 0.6 389.0 $\pm$ 1.5 173.0 $\pm$ 2.7 2159.2 $\pm$ 48.4 506.2 $\pm$ 2.6
Oracle GP-TS 35.0 $\pm$ 1.1 28.1 $\pm$ 0.8 86.0 $\pm$ 1.0
Oracle GP-UCB 68.5 $\pm$ 1.9 48.3 $\pm$ 1.2 217.3 $\pm$ 1.0
SCoreBO 180.4 $\pm$ 7.7 180.8 $\pm$ 5.8 106.6 $\pm$ 0.9 256.8 $\pm$ 9.3 3460.2 $\pm$ 163.7 861.3 $\pm$ 21.0
EEI 39.0 $\pm$ 2.6 30.1 $\pm$ 2.1 88.3 $\pm$ 4.2 51.6 $\pm$ 4.8 1664.1 $\pm$ 137.7 196.5 $\pm$ 12.3
First, we provide the average total regret for the synthetic and real-world data experiments in Table˜1. We observe that HP-GP-TS either has the lowest regret or is within 1 standard error of the algorithm with the lowest regret across all the experiments. In Figure˜3, it can be noted had low regret early in the synthetic experiment but HP-GP-TS either catches up or almost catches up later in the experiments. We compare both algorithms with an extended time horizon of $T=1500$, the results are shown in Figures˜10 and 2. With the extended time horizon, HP-GP-TS achieves the lowest regret across all synthetic experiments. Although, EEI is still within 1 standard error on the lengthscale experiment.
Figure 10: Cumulative regret for synthetic experiments extended time horizon $T=1500$ with varying kernel (left), lengthscale (center) and mean function (right). Errorbars correspond to $\pm 1$ standard error.
Table 2: Average total regret and $\pm$ 1 standard error for the synthetic experiments with longer horizon $T=1500$. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Kernel Lengthscale Subspace
HP-GP-TS 49.1 $\pm$ 1.6 39.7 $\pm$ 1.2 103.4 $\pm$ 1.3
EEI 62.6 $\pm$ 6.2 45.7 $\pm$ 5.0 138.9 $\pm$ 9.2
Figure 11: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
Figure 12: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS. The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
Next, we include the mean number of priors in $P_{t}$ for all experiments in Figure˜11. Similarly, we include the average entropy of the hyperposterior for all experiments in Figure˜12. For the lengthscale, subspace, PeMS and PNW precipitation experiments, hardly any priors are eliminated. In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipitation having the most and least concentrated hyperposteriors.
The cost of learning the prior, ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$, for HP-GP-TS is shown in Figure˜13. Whilst this term has not been bounded theoretically in Theorem˜4.2, we observe that empirically its growth diminishes much quicker than the cumulative regret. Since ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is one term in the upper bound for the regret of HP-GP-TS, learning the prior does not seem to increase the growth rate of the regret empirically.
Figure 13: The cost of learning the prior, ${\sum_{t\in[T]}}U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})$, for HP-GP-TS in the synthetic experiments. The solid blue line corresponds to the mean and the dashed lines corresponds to the first and last decile at each time step.
We include the full set of confusion matrices for the lengthscale and subspace experiments in Figure˜14. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively. Even though PE-GP-UCB has a higher accuracy than PE-GP-TS, it still has significantly higher regret. Additionally, the priors are equivalent up to coordinate permutations and therefore generate functions that are equally difficult to optimize. Unlike the kernel and lengthscale experiments, the PE-methods do not oversample any specific prior but commit too much time to exploring along the irrelevant dimensions.
In Tables˜3 and 4, the total regret for the lengthscale and subspace scaling experiments are shown.
(a) Kernel experiment
(b) Lengthscale experiment
(c) Subspace experiment
Figure 14: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic experiments.
Table 3: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Lengthscales, $|P|$
8 16 32 64 128
MAP GP-TS 30.2 $\pm$ 1.2 32.4 $\pm$ 2.5 32.5 $\pm$ 2.1 28.7 $\pm$ 1.1 30.8 $\pm$ 1.9
HP-GP-TS 31.4 $\pm$ 1.0 31.7 $\pm$ 0.9 30.8 $\pm$ 0.8 30.7 $\pm$ 1.0 31.0 $\pm$ 1.4
PE-GP-TS 61.8 $\pm$ 0.5 61.3 $\pm$ 0.5 62.2 $\pm$ 0.5 62.4 $\pm$ 0.4 64.3 $\pm$ 0.4
PE-GP-UCB 114.2 $\pm$ 0.6 114.8 $\pm$ 0.6 115.5 $\pm$ 0.6 114.5 $\pm$ 0.6 114.8 $\pm$ 0.6
Oracle GP-TS 28.1 $\pm$ 0.8 26.4 $\pm$ 0.8 27.3 $\pm$ 0.8 26.5 $\pm$ 0.7 25.7 $\pm$ 0.7
Oracle GP-UCB 48.3 $\pm$ 1.2 46.9 $\pm$ 1.1 48.4 $\pm$ 1.1 46.5 $\pm$ 1.0 45.6 $\pm$ 1.0
SCoreBO 180.8 $\pm$ 5.8 240.3 $\pm$ 6.3 277.2 $\pm$ 6.8 281.6 $\pm$ 6.7 283.9 $\pm$ 6.8
EEI 30.1 $\pm$ 2.1 30.9 $\pm$ 2.2 29.4 $\pm$ 2.2 30.8 $\pm$ 2.6 32.1 $\pm$ 2.8
Table 4: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Subspaces, $|P|$
5 8 12 16
MAP GP-TS 87.2 $\pm$ 1.0 89.9 $\pm$ 1.1 89.1 $\pm$ 0.9 90.9 $\pm$ 1.2
HP-GP-TS 88.3 $\pm$ 0.9 88.8 $\pm$ 0.9 89.5 $\pm$ 0.9 90.8 $\pm$ 0.9
PE-GP-TS 177.1 $\pm$ 1.4 269.5 $\pm$ 1.9 344.7 $\pm$ 2.3 396.9 $\pm$ 2.5
PE-GP-UCB 389.0 $\pm$ 1.5 526.0 $\pm$ 1.8 622.4 $\pm$ 2.3 688.0 $\pm$ 2.7
Oracle GP-TS 86.0 $\pm$ 1.0 84.1 $\pm$ 0.9 84.6 $\pm$ 1.0 84.8 $\pm$ 1.0
Oracle GP-UCB 217.3 $\pm$ 1.0 218.2 $\pm$ 1.0 218.6 $\pm$ 1.0 218.9 $\pm$ 0.9
SCoreBO 106.6 $\pm$ 0.9 108.2 $\pm$ 0.8 108.9 $\pm$ 0.7 109.5 $\pm$ 0.7
EEI 88.3 $\pm$ 4.2 81.3 $\pm$ 3.8 82.5 $\pm$ 3.6 81.3 $\pm$ 3.8